OpenStudy (anonymous):

What is the value of the expression 8–2? A.–16 B. 1/64 C. 1/16 D.64

OpenStudy (just_one_last_goodbye):

Hey you should post this in math ^_^

OpenStudy (anonymous):

@mathmath333 @Tazmaniadevil

OpenStudy (just_one_last_goodbye):

@MadisonEmery is that 8 take away 2?

OpenStudy (anonymous):

ya

OpenStudy (just_one_last_goodbye):

give me a sec

OpenStudy (just_one_last_goodbye):

I really dont see how the answer choices are relevant really sorry im no help ;-;

OpenStudy (just_one_last_goodbye):

@pooja195 please help my friend ^_^

pooja195 (pooja195):

If this is a square 8^-2 then all you need to do is -8 x -8

OpenStudy (anonymous):

thx

pooja195 (pooja195):

:)

OpenStudy (anonymous):

so it would be D

OpenStudy (just_one_last_goodbye):

oh I didn't see the exponent

OpenStudy (just_one_last_goodbye):

thats why i asked if this was a subtraction problem not \[8^{-2}\]

OpenStudy (anonymous):

so it would be D @pooja195

OpenStudy (just_one_last_goodbye):

Yes @MadisonEmery a negative times a negative is a positive and 8 x 8 is 64

OpenStudy (anonymous):

this is wrong

OpenStudy (anonymous):

!!!!!@pooja195 !!!!!! you are wrong thats not how it works

pooja195 (pooja195):

?

OpenStudy (anonymous):

this should be: 1/8^-2

OpenStudy (anonymous):

therefore 1/64

OpenStudy (anonymous):

This is not multiplication, this is an exponent. With negative exponents you put the 8^-2 under 1 like this: 1/8^-2 the answer is NOT 64 the answer is 1/64

OpenStudy (anonymous):

Please look

pooja195 (pooja195):

well I was taught diffrently.

OpenStudy (just_one_last_goodbye):

Im afraid Twistnflip is correct. 8^-2 = 1/ 8 * 8 = 1/8^2 = 1/64

OpenStudy (just_one_last_goodbye):

sorry Pooja ;-; nothing personal

pooja195 (pooja195):

Also the problem was written wrongly.

OpenStudy (just_one_last_goodbye):

^ very

pooja195 (pooja195):

But i accept my mistake :)

OpenStudy (anonymous):

its ok, just wanted to make sure she didn't get it wrong

OpenStudy (just_one_last_goodbye):

Shes so mature ^-^